Review of ⟨Title of Talk⟩ Presented by ⟨Name of Speaker⟩ ⟨Speaker′s Affiliation⟩ as a Departmental Colloquium ⟨Date of Talk⟩

⟨Name⟩

Department of Computer Science and Engineering State University of New York at Buffalo 226 Bell Hall Buffalo, NY 14260-2000

                           ⟨login⟩@cs.buffalo.edu

⟨date of review⟩

1 Summary

Give a short summary of the main points of the talk in your own words, to the best of your ability.

2 Critique of Contents

Briefly describe your reaction to/impression of the contents of the talk.

3 Critique of Style

3.1 Positive Aspects

Briefly describe positive aspects of the speaker’s speaking/presentation style.

3.2 Negative Aspects

Briefly describe negative aspects of the speaker’s speaking/presentation style.